OpenStudy (anonymous):

How to find the sum of the masses of nh4cl, nacl and sio2?

OpenStudy (anonymous):

Add the atomic masses of each element in the compound. For example Mass of NH4Cl= Mass of N + 4*Mass of H + Mass of Cl = 14+4+35.5= Ans. Similarily find the masses of others and then add.